# Extreme Compression of Large Language Models via Additive Quantization

### 11 Jan 2024 v1

@@ Abstract
The emergence of accurate open large language models (LLMs) has led to a race towards quantization techniques for such models enabling execution on end-user devices. In this paper, we revisit the problem of “extreme” LLM compression–defined as targeting extremely low bit counts, such as 2 to 3 bits per parameter, from the point of view of classic methods in Multi-Codebook Quantization (MCQ). Our work builds on top of Additive Quantization, a classic algorithm from the MCQ family, and adapts it to the quantization of language models. The resulting algorithm advances the state-of-the-art in LLM compression, outperforming all recently-proposed techniques in terms of accuracy at a given compression budget. For instance, when compressing Llama 2 models to 2 bits per parameter, our algorithm quantizes the 7B model to 6.93 perplexity (a 1.29 improvement relative to the best prior work, and 1.81 points from FP16), the 13B model to 5.70 perplexity (a .36 improvement) and the 70B model to 3.94 perplexity (a .22 improvement) on WikiText2. We release our implementation of Additive Quantization for Language Models AQLM as a baseline to facilitate future research in LLM quantization.
@@ 1 Introduction
(2-bit) and the original 16-bit weights on Llama 2 models.
The rapid advancement of generative large language models (LLMs) has led to massive industrial and popular interest, driven in part by the availability of accurate open LLMs, such as Llama 1 and 2 (Touvron et al., 2023), Falcon (TII UAE, 2023), BLOOM (Scao et al., 2022), OPT (Zhang et al., 2022), or NeoX/Pythia (Biderman et al., 2023).
@@
A key advantage of such open models is that they can be inferenced or fine-tuned locally by end-users, assuming that their computational and memory costs can be brought down to be manageable on commodity hardware. Consequently, this has led to interest in methods for inference and fine-tuning on compressed LLMs (Dettmers et al., 2022; Frantar et al., 2022a; Dettmers & Zettlemoyer, 2022; Lin et al., 2023; Dettmers et al., 2023a). Currently, the primary approach for accurate post-training compression of LLMs is quantization, which reduces the bit-width at which model weights (and possibly activations) are stored, leading to improvements in model footprint and memory transfer.
@@
By and large, the general approach to LLM weight compression can be described as “direct” quantization, in the sense that a suitable quantization grid and normalization are first chosen for each matrix sub-component, and then weights are mapped onto the grid either by direct rounding, e.g. (Dettmers & Zettlemoyer, 2022), or via more complex allocations, e.g. (Frantar et al., 2022a). Quantization induces a natural compression-vs-accuracy trade-off, usually measured in terms of bits-per-element versus model perplexity (PPL). State-of-the-art approaches can achieve arguably low accuracy loss at 3-4 bits per element (Dettmers et al., 2023b; Chee et al., 2023; Kim et al., 2023), and can even stably compress models to 2 bits per element. Yet, with current methods, low bit counts come at the cost of significant drops in accuracy, higher implementation complexity and runtime overheads. Specifically, from a practitioner’s perspective, “extreme” quantization in the 2-bit range using current techniques is inferior to simply using a smaller base model and quantizing it to higher bitwidths, such as 3-4 bits per parameter, as the latter approach usually yields higher accuracy given the same model size in bytes (Dettmers & Zettlemoyer, 2022; Chee et al., 2023).
@@
In this work, we aim to improve the state-of-the-art in the high-compression range. We investigate an alternative approach to LLM compression, by extending Multi-Codebook Quantization (MCQ) to LLMs: broadly, MCQ is a family of methods developed in the context of information retrieval and approximate nearest-neighbor (ANN) search (Chen et al., 2010; Jegou et al., 2010; Ge et al., 2013; Zhang et al., 2014; Babenko & Lempitsky, 2014; Martinez et al., 2016, 2018), consisting of specialized quantization algorithms to compress databases of vectors, allowing for efficient search against this database. Unlike direct quantization algorithms, vector quantization compresses multiple values together, by leveraging the mutual information of quantized values.
@@
Throughout this paper, we extend vector quantization from approximate nearest neighbor search to the task of compressing LLM weights with low compression error. More specifically, we adapt Additive Quantization (AQ) (Babenko & Lempitsky, 2014; Martinez et al., 2016) — a popular vector quantization algorithm — to the task of compressing LLM weights such that the output of each layer and Transformer block is approximately preserved. To that end, we reformulate the AQ optimization problem to reduce the error in LLM layer outputs, rather than preserving the weights themselves. We refer to the resulting procedure as Additive Quantization of Language Models (AQLM). Unlike other extreme LLM quantization approaches (Kim et al., 2023; Dettmers et al., 2023b), AQ quantizes models in a simple homogeneous format, which should be easy to implement in practice. The main contributions of this work can be summarized as follows:
@@
    We propose a practical adaptation of Additive Quantization to the task of post-training quantization of LLMs. Our technical contribution is two-fold: (1) we adapt the underlying optimization problem (specifically, a Map MRF) to be instance-aware, taking layer calibration input & output activations into account; (2) we complement the layer-wise optimization with an efficient intra-layer tuning technique, which optimizes quantization parameters jointly over several layers, using only the calibration data.
@@
    We evaluate the effectiveness of this algorithm on the task of compressing 7-70B parameter LLMs from the Llama2 family (Touvron et al., 2023) with compression rates of 2-4 bits per parameter. We find that AQLM outperforms the previous state-of-the-art algorithm across the 2-4 bit compression, with most significant improvements for extreme 2-bit quantization.
@@
    To facilitate reproducibility, we publish the reference implementation [^1] https://github.com/vahe1994/AQLM of AQLM, along with several pre-quantized models.
@@ 2.1 LLM Quantization
We first overview work on post-training quantization (PTQ) methods (Nagel et al., 2020; Gholami et al., 2021) in the context of LLMs. Broadly, while there is significant work on classic PTQ techniques such as AdaRound (Nagel et al., 2020), BitSplit (Wang et al., 2020), AdaQuant (Hubara et al., 2021), BRECQ (Li et al., 2021), and OBQ (Frantar et al., 2022b); these methods primarily focus on accurate quantization of smaller models, e.g. in the 10-100M range, and are hard to scale for LLMs due to the computational demands of their underlying solvers for billion-scale models.
@@
Early efforts towards PTQ methods that scale to LLMs such as ZeroQuant (Yao et al., 2022), LLM.int8() (Dettmers et al., 2022), and nuQmm (Park et al., 2022) employed direct round-to-nearest (RTN) projections, and adjusted quantization granularity to balance memory efficiency and accuracy. GPTQ (Frantar et al., 2022a) proposed a more accurate data-aware approach by leveraging an approximate large-scale solver for minimizing layer-wise $\ell_{2}$ errors.
@@
Dettmers et al. (Dettmers & Zettlemoyer, 2022) examined the accuracy-compression trade-offs of these early methods, suggesting that 4-bit quantization may be optimal for RTN quantization, and observing that data-aware methods like GPTQ allow for higher compression, i.e. strictly below 4 bits per weight, still maintaining Pareto optimality. Parallel work quantizing both weights and activations to 8-bits, by Dettmers et al. (2022)Xiao et al. (2022), and Yao et al. (2022) yielded insights into the sources of quantization error, notably the fact that the “outlier features” in large LLMs cause substantial quantization errors, prompting various mitigation strategies.
@@
Following this initial wave of work on LLM quantization, several improved techniques have been proposed, mainly focusing on the difficulty of quantizing outliers in the weight matrix. SpQR (Dettmers et al., 2023b) addresses this by saving the outlier weights—which have disproportionately high impact on the output error—as a highly-sparse higher-precision matrix. AWQ (Lin et al., 2023) reduces the error of quantizing channels with the highest activation magnitudes by employing per-channel scaling to reduce the error on important weights. SqueezeLLM (Kim et al., 2023) uses the diagonal Fisher as a proxy for the Hessian and implements non-uniform quantization through K-means clustering. This method identifies outliers as large entries, or entries sensitive to quantization, retaining them in original precision.
Currently, the state-of-the-art method in terms of accuracy-to-size trade-off is QuIP (Chee et al., 2023), and its newer variant QuIP# (Tseng et al., ). Roughly, these methods work by first “smoothening” outliers by multiplying with a rotation matrix (e.g., a Hadamard transform), and then mapping the smoothened weights onto a lattice via a projection procedure. QuIP was the first method to obtain stable results (i.e., single-digit PPL increases) in the 2-bit per parameter compression range.
@@
At a high level, these methods aim to minimize the “worst-case” quantization error for each layer, given initial weights and calibration data. For instance, in QuIP#, the distribution of the rotated weights approximates a Gaussian distribution, while the encoding lattice (E8P) is chosen to minimize the “rounding” error. By contrast, our approach uses a different weight encoding (codebooks are used additively), and learned codebooks instead of a fixed codebook, leveraging the intuition that we should be able to obtain better encoding by direct optimization. In the same vein, we allow codebooks for different layers to co-train via a separate fine-tuning procedure.
@@ 2.2 Quantization for Nearest Neighbor Search
Our investigation builds on a different family of quantization algorithms designed for approximate nearest neighbor search (ANN). Unlike PTQ, the main objective of ANN quantization is to compress a database of vectors in a way that allows user to efficiently compute similarities and find nearest neighbors, typically in terms of euclidean distance or inner product, relative to a set of query points. To achieve high compression rates, modern ANN search algorithms can use vector quantization (VQ) — a technique that quantizes multiple vector dimensions together, as a vector (Burton et al., 1983; Gray, 1984). It achieves this by learning “codebooks”: a codebook is a set of learnable candidate vectors that can be used to encode the data.
To encode a given database vector, VQ splits it into sub-groups of several dimensions each, then encodes every group by choosing a vector from the learned codebook. The algorithm was designed to efficiently compute distances or dot-products for similarity search by taking advantage of the linear properties of dot products.
@@
Modern quantization methods for ANN search generalize the idea of vector quantization and are referred to as multi-codebook quantization (MCQ) in the literature. MCQ methods typically do not involve the information loss on the query side, what makes them the leading paradigm for the memory-efficient ANN (Ozan et al., 2016; Martinez et al., 2018). We review several MCQ variants below.
@@
Product quantization (PQ) (Jegou et al., 2010) is a pioneering method from the MCQ family, which encodes each vector $x\in\mathbf{R}^{D}$ as a concatenation of $M$ codewords from $M$ $\frac{D}{M}$-dimensional codebooks $C_{1},\ldots,C_{M}$, each containing $K$ codewords. Simply put, PQ decomposes a vector into $M$ separate subvectors and applies vector quantization (VQ) to each subvector, while using a separate codebook. As a result, each vector $x$ is encoded by a tuple of codeword indices $[i_{1},\ldots,i_{M}]$ and approximated by $x\approx[c_{1i_{1}},\ldots,c_{Mi_{M}}]$. Fast Euclidean distance computation becomes possible using lookup tables:
$\|q-x\|^{2}\approx\|q-[c_{1i_{1}},\ldots,c_{Mi_{M}}]\|^{2}=$ (1)
$\sum\limits_{m=1}^{M}{\|q_{m}-c_{mi_{m}}\|}^{2}$
where $q_{m}$ is the $m$th subvector of a query $q$. This sum can be calculated using $M$ additions and lookups if the distances from query subvectors to codewords are precomputed.
Geometrically, PQ effectively partitions the original vector space into $K^{M}$ “cells,” each being a Cartesian product of $M$ lower-dimensional cells. Such product-based approximation works better if the $\frac{D}{M}$-dimensional components of vectors have independent distributions. The degree of dependence is affected by the choice of the splitting, and can be further improved by orthogonal transformation applied to vectors as pre-processing. Two subsequent works have therefore looked into finding an optimal transformation (Ge et al., 2013; Norouzi & Fleet, 2013). Modifications of PQ corresponding to such pre-processing transformations are referred below as Optimized Product Quantization (OPQ).
@@
Non-orthogonal quantizations. Follow-up works (Chen et al., 2010; Babenko & Lempitsky, 2014; Martinez et al., 2016; Zhang et al., 2014; Ozan et al., 2016; Martinez et al., 2018) generalized the idea of Product Quantization by approximating each vector by a sum of $M$ codewords instead of concatenation. The resulting procedure is still efficient while the approximation accuracy is increased.
@@
The first approach, Residual Vector Quantization (Chen et al., 2010), quantizes original vectors, and then iteratively quantizes the approximation residuals from the previous iteration. Another approach, Additive Quantization (AQ) (Babenko & Lempitsky, 2014), is the most general as it does not impose any constraints on the codewords from the different codebooks. Usually, AQ provides the smallest compression errors, however, it is much slower than other methods, especially for large $M$. This is because AQ relies on an iterative two-phase optimization procedure that alternates between updating codebooks and re-encoding the database. Since this procedure is central to our work, we discuss it in more detail in Section 3.
@@
Finally, several recent works (Martinez et al., 2016, 2018; Zhang et al., 2014) elaborate the idea of Additive Quantization, proposing the more effective procedure for codebooks learning. Composite Quantization (CQ) (Zhang et al., 2014) learns codebooks with a fixed value of scalar product between the codewords from different codebooks. Currently, state-of-the-art compression accuracy is achieved by the LSQ method (Martinez et al., 2018).
@@
There has been significant work on exploiting vector quantization in the context of machine learning. For instance, Zhou et al. (2017); Li et al. (2017); Chen et al. (2019) use multi-codebook quantization to compress word embeddings within deep learning models. Another line of work (Blalock & Guttag, 2021; McCarter & Dronen, 2022; Fernández-Marqués et al., 2023) explores vector quantization for linear models, or linear layers within deep models. Similarly to PQ above, these techniques pre-compute inner products between inputs and all codes, then compute linear layer via look-up, which speeds up inference. However, these algorithms introduce significant prediction error that does not allow them to compress deep models. In this context, we are the first to successfully adapt MCQ approaches to LLMs.

### 06 Feb 2024 v2

@@ Abstract
The emergence of accurate open large language models (LLMs) has led to a race towards performant quantization techniques which can enable their execution on end-user devices. In this paper, we revisit the problem of “extreme” LLM compression—defined as targeting extremely low bit counts, such as 2 to 3 bits per parameter—from the point of view of classic methods in Multi-Codebook Quantization (MCQ). Our algorithm, called AQLM, generalizes the classic Additive Quantization (AQ) approach for information retrieval to advance the state-of-the-art in LLM compression, via two innovations: 1) learned additive quantization of weight matrices in input-adaptive fashion, and 2) joint optimization of codebook parameters across entire layer blocks. Broadly, AQLM is the first scheme that is Pareto optimal in terms of accuracy-vs-model-size when compressing to less than 3 bits per parameter, and significantly improves upon all known schemes in the extreme compression (2bit) regime. In addition, AQLM is practical: we provide fast GPU and CPU implementations of AQLM for token generation, which enable us to match or outperform optimized FP16 implementations for speed, while executing in a much smaller memory footprint.
@@ 1 Introduction
Figure 1: Comparison of AQLM (2-bit) relative to the state-of-the-art QuIP# (2-bit) and the original 16-bit weights on Llama 2 7, 13, and 70B models.
@@
The rapid advancement of generative large language models (LLMs) has led to massive industrial and popular interest, driven in part by the availability of accurate open LLMs, such as Llama 1 and 2 (Touvron et al., 2023), Falcon (TII UAE, 2023), BLOOM (Scao et al., 2022), OPT (Zhang et al., 2022), or NeoX/Pythia (Biderman et al., 2023). A key advantage of open models is that they can be inferenced or fine-tuned locally by end-users, assuming that their computational and memory costs can be reduced to be manageable on commodity hardware. This has led to several methods for inference and fine-tuning on compressed LLMs (Dettmers et al., 2022; Frantar et al., 2022a; Dettmers & Zettlemoyer, 2022; Lin et al., 2023; Dettmers et al., 2023a). Currently, the primary approach for accurate post-training compression of LLMs is quantization, which reduces the bit-width at which model weights (and possibly activations) are stored, leading to improvements in model footprint and memory transfer.
@@
By and large, LLM weights are compressed via “direct” quantization, in the sense that a suitable quantization grid and normalization are first chosen for each matrix sub-component, and then weights are each mapped onto the grid either by direct rounding, e.g. (Dettmers & Zettlemoyer, 2022), or via more complex allocations, e.g. (Frantar et al., 2022a). Quantization induces a natural compression-vs-accuracy trade-off, usually measured in terms of model size vs model perplexity (PPL). Existing approaches can achieve arguably low accuracy loss at 3-4 bits per element (Dettmers et al., 2023b; Chee et al., 2023; Kim et al., 2023), and can even stably compress models to 2 or even less bits per element, in particular, for extremely large models (Frantar & Alistarh, 2023). Yet, in most cases, low bit counts come at the cost of significant drops in accuracy, higher implementation complexity and runtime overheads. Specifically, from the practical perspective, “extreme” quantization in the 2-bit range using current techniques is inferior to simply using a smaller base model and quantizing it to higher bitwidths, such as 3-4 bits per parameter, as the latter yields higher accuracy given the same model size in bytes (Dettmers & Zettlemoyer, 2022; Chee et al., 2023).
@@
Contribution. In this work, we improve the state-of-the-art in LLM compression by showing for the first time that Multi-Codebook Quantization (MCQ) techniques can be extended to LLM weight compression. Broadly, MCQ is a family of information retrieval methods (Chen et al., 2010; Jegou et al., 2010; Ge et al., 2013; Zhang et al., 2014; Babenko & Lempitsky, 2014; Martinez et al., 2016, 2018), consisting of specialized quantization algorithms to compress databases of vectors, allowing for efficient search. Unlike direct quantization, MCQ compresses multiple values jointly, by leveraging the mutual information of quantized values.
@@
More precisely, we extend Additive Quantization (AQ) (Babenko & Lempitsky, 2014; Martinez et al., 2016), a popular MCQ algorithm, to the task of compressing LLM weights such that the output of each layer and Transformer block are approximately preserved. Our extension reformulates the classic AQ optimization problem to reduce the error in LLM layer outputs under the input token distribution and as well as to jointly optimize codes over layer blocks, rather than only preserving the weights themselves as in standard AQ. We refer to the resulting procedure as Additive Quantization of Language Models (AQLM). Unlike some extreme LLM quantization approaches that require hybrid sparse-quantized formats which separate outlier quantization (Kim et al., 2023; Dettmers et al., 2023b), AQLM quantizes models in a simple homogeneous format, which is easy to support in practice. Our main contributions are as follows:
@@
    We propose the AQLM algorithm, which extends AQ to post-training compression of LLM weights, via two innovations: (1) adapting the MAP-MRF optimization problem behind AQ to be instance-aware, taking layer calibration input & output activations into account; (2) complementing the layer-wise optimization with an efficient intra-layer tuning technique, which optimizes quantization parameters jointly over several layers, using only the calibration data.
@@
    We evaluate the effectiveness of this algorithm on the task of compressing accurate open LLMs from the Llama 2 (Touvron et al., 2023) family with compression rates of 2-4 bits per parameter. We find that AQLM outperforms the previous state-of-the-art across the standard 2-4 bit compression range, with the most significant improvements for extreme 2-bit quantization (see Figure 1). We provide detailed ablations for the impact of various algorithm parameters, such as code width and number of codebooks, and extend our analysis to the recent Mixtral model (Jiang et al., 2024).
@@
    We show that AQLM is practical, by providing efficient GPU and CPU kernels implementations for specific encodings, as well as end-to-end generation [^1] https://github.com/vahe1994/AQLM. Results show show that our approach can match or even outperform the floating point baseline in terms of speed, while reducing the memory footprint by up to 8x. Specifically, AQLM can be executed with layer-wise speedups of $\sim 30\%$ for GPUs, and of up to 4x for CPU inference.
@@ 2.1 LLM Quantization
Early efforts towards post-training quantization (PTQ) methods (Nagel et al., 2020; Gholami et al., 2021) that scale to LLMs such as ZeroQuant (Yao et al., 2022), LLM.int8() (Dettmers et al., 2022), and nuQmm (Park et al., 2022) employed direct round-to-nearest (RTN) projections, and adjusted quantization granularity to balance memory efficiency and accuracy. GPTQ (Frantar et al., 2022a) proposed a more accurate data-aware approach via an approximate large-scale solver for minimizing layer-wise $\ell_{2}$ errors.
@@
Dettmers & Zettlemoyer (2022) examined the accuracy-compression trade-offs of these early methods, suggesting that 4-bit quantization may be optimal for RTN quantization, and observing that data-aware methods like GPTQ allow for higher compression, i.e. strictly below 4 bits/weight, maintaining Pareto optimality. Our work brings this Pareto frontier below 3 bits/weight, for the first time. Parallel work quantizing both weights and activations to 8-bits, by Dettmers et al. (2022)Xiao et al. (2022), and Yao et al. (2022) noted that the “outlier features” in large LLMs cause substantial errors, prompting various mitigation strategies.
@@
Recently, several improved techniques have focused on the difficulty of quantizing weight outliers, which have high impact on the output error. SpQR (Dettmers et al., 2023b) addresses this by saving outliers as a highly-sparse higher-precision matrix. AWQ (Lin et al., 2023) reduces the error of quantizing channels with the highest activation magnitudes by employing per-channel scaling to reduce the error on important weights. SqueezeLLM (Kim et al., 2023) uses the diagonal Fisher as a proxy for the Hessian and implements non-uniform quantization through K-means clustering.
@@
The state-of-the-art method in terms of accuracy-to-size trade-off is QuIP (Chee et al., 2023). Concurrent to our work, an improved variant called QuIP# (Tseng et al., 2023) was introduced. Roughly, these methods work by first “smoothening” weights by multiplying with a rotation matrix, and then mapping them onto a lattice. QuIP was the first method to obtain stable results (i.e., single-digit PPL increases) in the 2-bit per parameter compression range. At a high level, QuIP and QuIP# aim to minimize the “worst-case” error for each layer, given initial weights and calibration data. For instance, in QuIP#, the distribution of the rotated weights approximates a Gaussian, while the encoding lattice (E8P) is chosen to minimize “rounding” error. By contrast, our approach uses a different weight encoding (codebooks are additive), and learned codebooks instead of a fixed codebook. Thus, our insight is that we should be able to obtain higher accuracy by direct optimization of the codebooks over the calibration set, removing the rotation. Further, we show that codebooks for different layers can co-train via joint fine-tuning over the calibration data.
@@ 2.2 Quantization for Nearest Neighbor Search
Our work builds on approximate nearest neighbor search (ANN) algorithms. Unlike PTQ, ANN quantization aims to compress a database of vectors to allow a user to efficiently compute similarities and find nearest neighbors relative to a set of query points. For high compression, modern ANN search algorithms employ vector quantization (VQ)—which quantizes multiple vector dimensions jointly (Burton et al., 1983; Gray, 1984). It achieves this by learning “codebooks”: i.e. a set of learnable candidate vectors that can be used to encode the data. To encode a given database vector, VQ splits it into sub-groups of entries, then encodes every group by choosing a vector from the learned codebook. The algorithm efficiently computes distances or dot-products for similarity search by leveraging the linearity of dot products.
@@
Quantization methods for ANN search generalize vector quantization and are referred to as multi-codebook quantization (MCQ). MCQ methods typically do not involve information loss on the query side, which makes them the leading approach for memory-efficient ANN (Ozan et al., 2016; Martinez et al., 2018). We briefly review MCQ below.
@@
Product quantization (PQ) (Jegou et al., 2010) is an early version of MCQ, which encodes each vector $x\in\mathbf{R}^{D}$ as a concatenation of $M$ codewords from $M$ $\frac{D}{M}$-dimensional codebooks $C_{1},\ldots,C_{M}$, each containing $K$ codewords. PQ decomposes a vector into $M$ separate subvectors and applies vector quantization (VQ) to each subvector, while using a separate codebook. Thus, each vector $x$ is encoded by a tuple of codeword indices $[i_{1},\ldots,i_{M}]$ and approximated by $x\approx[c_{1i_{1}},\ldots,c_{Mi_{M}}]$. Fast Euclidean distance computation becomes possible using lookup tables:
$\|q-x\|^{2}\approx\|q-[c_{1i_{1}},\ldots,c_{Mi_{M}}]\|^{2}=\sum\limits_{m=1}^{M}{\|q_{m}-c_{mi_{m}}\|}^{2},$
where $q_{m}$ is the $m$th subvector of a query $q$. This sum can be calculated using $M$ additions and lookups if the distances from query subvectors to codewords are precomputed. Since product-based approximations work better if the $\frac{D}{M}$-dimensional components independent distributions, subsequent work has looked into finding better transformations (Ge et al., 2013; Norouzi & Fleet, 2013). As for the other similarity functions, (Guo et al., 2016) proposes a quantization procedure for maximum inner product search (MIPS). They minimize quantization error in the inner products between database and query vectors by solving a constrained optimization problem. Similarly to the formula above, this procedure allows for efficient inner product search by precomputing dot products between the query $q$ an all codes in the learned codebooks, then adding these partial dot products to recover the full similarity score.
@@
Non-orthogonal quantizations. Follow-up work (Chen et al., 2010; Babenko & Lempitsky, 2014; Martinez et al., 2016; Zhang et al., 2014; Ozan et al., 2016; Martinez et al., 2018) generalized the idea of Product Quantization by approximating each vector by a sum of $M$ codewords instead of concatenation. The resulting procedure is still efficient while the approximation accuracy is increased.
@@
For this, Residual Vector Quantization (Chen et al., 2010), quantizes original vectors, and then iteratively quantizes the approximation residuals from the previous iteration. Additive Quantization (AQ) (Babenko & Lempitsky, 2014) is more general, as it does not impose constraints on the codewords from the different codebooks. Usually, AQ provides the smallest compression errors, but is more complex to train for large $M$. We discuss this in detail in Section 3.
@@
Finally, several recent works (Martinez et al., 2016, 2018; Zhang et al., 2014) elaborate the idea of Additive Quantization, proposing the more effective procedure for codebooks learning. Composite Quantization (CQ) (Zhang et al., 2014) learns codebooks with a fixed value of scalar product between the codewords from different codebooks. Currently, the state-of-the-art compression accuracy is achieved by the LSQ method (Martinez et al., 2018).
@@
There has been significant work on exploiting vector quantization in the context of machine learning. For instance, Zhou et al. (2017); Li et al. (2017); Chen et al. (2019) use multi-codebook quantization to compress word embeddings within deep learning models. Another line of work (Blalock & Guttag, 2021; McCarter & Dronen, 2022; Fernández-Marqués et al., 2023) explores vector quantization for linear models, or linear layers within deep models. Similarly to PQ above, these techniques pre-compute inner products between inputs and all codes, then compute linear layer via look-up, which speeds up inference. However, these algorithms introduce significant prediction error that does not allow them to compress deep models. Thus, we believe we are the first to successfully adapt and scale MCQ to LLMs.